OpenStudy (anonymous):

If f(x) = 1/2x-3 and g(x)= 2x+5 what is (g x f) (4)

OpenStudy (anonymous):

\[ \begin{split} (g\times f)(4) &= g(4) \times f(4) \\ &= [2(4)+5] \times [1/2(4)-3] \end{split} \]Think you can do the rest?

OpenStudy (anonymous):

Also, it matters whether \(f(x) = \frac{1}{2}x-3\) or \(f(x) = \frac{1}{2x}-3\) or even \( f(x) = \frac{1}{2x-3} \)

OpenStudy (anonymous):

Thanks so much this was very helpful :) !

OpenStudy (anonymous):

Is the answer -13 ?

OpenStudy (anonymous):

It depends on which f(x) is correct.